0-6[2-12(2-5)+1/2(20)-42+5-3]+0-10
1
Expert's answer
2019-09-25T09:48:58-0400

You have the following mathematical operation


"0-6[2-12(2-5)+\frac{1}{2}(20)-42+5-3]+0-10"


We perform the operations between parentheses


"0-6[2-12(-3)+\frac{20}{2}-42+5-3]+0-10\\ 0-6[2+36+10-42+5-3]+0-10"



Now the algebraic sum is made

"2+36+10-42+5-3=8"

"0-10=-10"


Now


"-6[8]-10" -->"-6[8]=-48"


Continuing


"-48-10=-58"


FInally


Solution "0-6[2-12(2-5)+\frac{1}{2}(20)-42+5-3]+0-10=-58"